OpenStudy (anonymous):

PLEASE HELP ASAP I DONT GET IT O.o find the value of y for a given value of x, if y varies directly with x if y=6 x=2, what is y when x=5 explain it in steps

OpenStudy (anonymous):

6/2 = y/5 y = 15

OpenStudy (the_fizicx99):

If y varies directly to x then 6/2 = 3 I think that if x = 5 then y would be equal to 15 What do you think @ranga

OpenStudy (ranga):

Yes. 15.

OpenStudy (the_fizicx99):

@rockstargirl17 do you understand what I tried to explain?

OpenStudy (anonymous):

yes and no so wait that means the answer is 15 im confused

OpenStudy (the_fizicx99):

Well see because y varies directly to x, that means that if you divide 6/2 it is equal to 3 so if you were trying to find y when x = 5 you multiply by 3 5 * 3 = 15 So yes the answer is 15 :)

OpenStudy (anonymous):

oh okay i get it thank you very much :) could you help me with another problem im stuck on this one too :P

OpenStudy (the_fizicx99):

Sure

OpenStudy (anonymous):

okay so it says Divid -2x^3+21x^2-20x-12 by (x-3)

OpenStudy (anonymous):

how exactly are you supose to do that :P

OpenStudy (the_fizicx99):

Divide?

OpenStudy (anonymous):

yeah i know that but how would you divid it like do you just combine like terms or what

OpenStudy (the_fizicx99):

What grade is this for? I wouldn't combine like terms because their exponents are different.

OpenStudy (anonymous):

11th

OpenStudy (the_fizicx99):

9th :< I'd put the fraction over a common denominator. -2(x-3)x^3/x-3 + 21(x-3)x^2/x-3 - 20(x-3)x - 12/x-3

OpenStudy (anonymous):

:( hmm

OpenStudy (the_fizicx99):

Combine it into a single fraction.

OpenStudy (the_fizicx99):

Ok then? :/

OpenStudy (anonymous):

i still dont get it :P

OpenStudy (the_fizicx99):

You're suppose to solve it. Putting it into a single fraction then distribute -2x^3 over x-3. I made an attempt to try to help you. :/